Eric earns a base salary of $50,000 as a shipwreck
Answer #1

Marginal tax rate on dive will be 50% as the Eric already earns 50,000 , so any extra earnings above that amount will be under 50% tax bracket.

Eric will not take the dive as it will earn him less after tax dollar as against expectations. He will get after tax = 5000*(1-50%) = 2,500 , while he needs 3,000 to take dive.

Eric's average tax rate will be = (10000*0.05+20000*0.10+20000*0.20+5000*0.50)/ 55000

= (500+2000+4000+2500) / 55000 = 16.36%

Total taxes paid (including extra dive) = (10000*0.05+20000*0.10+20000*0.20+5000*0.50) = 9,000

Total taxes paid (excluding extra dive) = (10000*0.05+20000*0.10+20000*0.20) = 6500